OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 2x − 3y = 6.

OpenStudy (anonymous):

slope = 2/3 and y-intercept = −2 slope = −2/3 and y-intercept = −2 slope = 2/3 and y-intercept = 2 slope = −2/3 and y-intercept = 2

OpenStudy (anonymous):

well what i would do is put it in y-intercept form

OpenStudy (anonymous):

get the -3y by itself...

OpenStudy (anonymous):

what would you do @usitdown

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

im sure if your in math you should know how to get the -3y by itself

OpenStudy (anonymous):

And im sure if i knew how to do this i wouldnt be asking the question.

OpenStudy (anonymous):

well if you REALLY dont know all you do is subtract 2x and get -3y=-2x+6 then what do you do to get y by itself? @usitdown

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

ok just dive the -3 and get y=2/3x-2

OpenStudy (anonymous):

thats your answer

OpenStudy (anonymous):

thanks